### v1

@@ Appendix 0.A Omitted proofs
To finish the proof, we argue that (de-)homogenization and cone polarity preserve extension complexity. The claim about (de-)homogenization follows from the discussion before [18, Theorem 6]. The statement about cone polarity is implied by the discussion before [18, Proposition 2]. ∎
Proof of Theorem 4.5.
@@
Consider the Newton polytope $P_{f}$ of the function $f$ computed by the neural network. Note that $P_{f}$ is contained in the nonnegative orthant since the weights of a monotone neural network are nonnegative. We slightly modify $P_{f}$ to also include all nonnegative vectors that are smaller than a vector in $P_{f}$ and obtain
$Q\coloneqq\{x\in\mathbb{R}^{E}\mid\exists y\in P_{f}\colon 0\leq x\leq y\}\ .$ (2)
@@
With that modification, $Q$ is a “monotone polytope” [^2] Note that this notion of a monotone polytope does not correspond to the notion of monotone neural networks. as defined in [37]. Moreover, we still maintain the inequality $f_{\mathrm{MAT}}(c)\leq f_{Q}(c)\leq(1+\epsilon)f_{\mathrm{MAT}}(c)$ for all $c\in\mathbb{R}_{\geq 0}^{E}$. As argued around [37, Equation 7], this is equivalent to $P_{\mathrm{MAT}}\subseteq Q\subseteq(1+\epsilon)P_{\mathrm{MAT}}$. By [37, Corollary 4.2], this implies that $\operatorname{xc}(Q)\geq 2^{\Omega(\min\{n,1/\epsilon\})}$; see also [39, 8]. The definition (2) of $Q$ implies $\operatorname{xc}(Q)\leq\operatorname{xc}(P_{f})+2\cdot\lvert E\rvert\leq\operatorname{xc}(P_{f})+\mathcal{O}(n^{2})$. Thus, we obtain the same asymptotic lower bound on $\operatorname{xc}(P_{f})$. The statement then follows by Theorem 4.1. ∎
Proof of Proposition 4.
As for $R$ the optimal face in $c$-direction is the singleton $\{x^{R}\}$, by Lemma 1 the corresponding faces of $P$ and $Q$ must be singletons, too, namely $\{x^{P}\}$ and $\{x^{Q}\}$ with $x^{P}+x^{Q}=x^{R}$. ∎
Proof of Proposition 5.
@@
For $i=1,\dots,d$, let $F^{P}_{i}$ be the set of optimal solutions over $P$ with respect to the lexicographic objective function $(c_{1}^{\top}x,\dots,c_{i}^{\top}x)$. Let $F^{Q}_{i}$ and $F^{R}_{i}$ be defined analogously. We show by induction on $i$ that $F^{P}_{i}+F^{Q}_{i}=F^{R}_{i}$. For $i=1$, this follows from Lemma 1. For $i>1$, observe that $F^{P}_{i}$ is the face in $c_{i}$-direction of $F^{P}_{i-1}$, and analogously for $Q$ and $R$. Hence, we can apply Lemma 1 again to complete the induction step. As a consequence, we obtain that $F^{P}_{d}+F^{Q}_{d}=F^{R}_{d}$. Moreover, since $C$ is a basis of $\mathbb{R}^{d}$, it follows that any two optimal solutions with respect to the lexicographic objective function must be equal, implying that $F^{P}_{d}$, $F^{Q}_{d}$, and $F^{R}_{d}$ are singletons. ∎
Proof of Theorem 5.1.
@@
Let $x^{P},x^{Q},x^{R}$ be the unique optimizers as guaranteed by Proposition 5. As they fulfill $x^{P}+x^{Q}=x^{R}$, the algorithm correctly returns $x^{P}=x^{R}-x^{Q}$. Furthermore, since $x^{P}$ is optimal with respect to the lexicographic objective, it is in particular optimal with respect to the first component of the objective, which is $c^{\top}x$. Therefore, the algorithm returns an optimal solution for $\max_{x\in P}c^{\top}x$.

### v2

@@ Appendix A Omitted Proofs
To finish the proof, we argue that (de-)homogenization and cone polarity preserve extension complexity. The claim about (de-)homogenization follows from the discussion before Gouveia et al. [2013, Theorem 6]. The statement about cone polarity is implied by the discussion before Gouveia et al. [2013, Proposition 2]. ∎
Proof of Theorem 3.9..
@@
Consider the Newton polytope $P_{f}$ of the function $f$ computed by the neural network. We slightly modify $P_{f}$ to obtain a polytope $Q$ in the following sense: firstly, we want to include only nonnegative vectors of $P_{f}$, and secondly, we also want to include all nonnegative vectors that are smaller than a vector in $P_{f}$. This can be formulated as
(2) $Q\coloneqq\{x\in\mathbb{R}^{E}\mid\exists y\in P_{f}\colon 0\leq x\leq y\}\ .$
With that modification, $Q$ is a “monotone polytope” [^4] Note that this notion of a monotone polytope does not correspond to the notion of monotone neural networks. as defined in Rothvoß [2017].
Let us analyze what we can say about the support function $f_{Q}$ of the modified polytope for $c\in\mathbb{R}_{\geq 0}^{E}$. Firstly, nonnegativity of $c$ implies $f_{Q}(c)\leq\max\{c^{\top}x\mid x\in P_{f}\cap\mathbb{R}_{\geq 0}^{E}\}\leq f(c)\leq(1+\epsilon)f_{\mathrm{MAT}}(c)$. Secondly, we claim that also $f_{Q}(c)\geq f_{\mathrm{MAT}}(c)$ holds for all $c\in\mathbb{R}_{\geq 0}^{E}$. To see this, assume the contrary, namely the existence of some $c\in\mathbb{R}_{\geq 0}^{E}$ with $f_{Q}(c)<f_{\mathrm{MAT}}(c)$. Let $x\in P_{\mathrm{MAT}}$ be an optimal vertex of the matching polytope in $c$-direction. It follows that $x\notin Q$. By definition of $Q$, it follows that $P_{f}$ and the set $X^{+}\coloneqq x+\mathbb{R}_{\geq 0}^{E}$ are disjoint. As these two sets are polyhedra, there must exist a separating hyperplane, that is, some vector $a\in\mathbb{R}^{E}$ and some value $\gamma\in\mathbb{R}$ with $a^{\top}y\leq\gamma$ for all $y\in P_{f}$ and $a^{\top}y>\gamma$ for all $y\in X^{+}$. The latter implies that $a$ is nonnegative, as the recession cone of $X^{+}$ is the nonnegative orthant. However, we then have $f_{\mathrm{MAT}}(a)\geq a^{\top}x>\gamma\geq f(a)$, contradicting the assumption of the theorem.
@@
In conclusion, also for $Q$ we have the inequality $f_{\mathrm{MAT}}(c)\leq f_{Q}(c)\leq(1+\epsilon)f_{\mathrm{MAT}}(c)$ for all $c\in\mathbb{R}_{\geq 0}^{E}$. As argued around Rothvoß [2017, Equation 7], this is equivalent to $P_{\mathrm{MAT}}\subseteq Q\subseteq(1+\epsilon)P_{\mathrm{MAT}}$. By Rothvoß [2017, Corollary 4.2], this implies that $\operatorname{xc}(Q)\geq 2^{\Omega(\min\{n,1/\epsilon\})}$; see also Sinha [2018], Braun and Pokutta [2014]. The definition (2) of $Q$ implies $\operatorname{xc}(Q)\leq\operatorname{xc}(P_{f})+2\cdot\lvert E\rvert\leq\operatorname{xc}(P_{f})+\mathcal{O}(n^{2})$. Thus, we obtain the same asymptotic lower bound on $\operatorname{xc}(P_{f})$. The statement then follows by Theorem 3.1. ∎
Proof of Proposition 4.2..
The proof is similar in spirit to Proposition 3.2, with some crucial differences making sure that we also eliminate negative weights between hidden neurons. In the end, this leads to two instead of one monotone network.
(3) $z_{v}$ $=\max\Big{\{}\sum_{u\in\delta^{\mathrm{in}}_{v}}\big{(}a_{uv}^{i}+\sum_{j\neq i}b_{uv}^{j}\big{)}z_{u}\Bigm{|}i=1,\dots,k\Big{\}}-\sum_{u\in\delta^{\mathrm{in}}_{v}}\sum_{j=1}^{k}b_{uv}^{j}z_{u}\enspace.$
Proof of Proposition 4.5..
@@
For $i=1,\dots,d$, let $F^{P}_{i}$ be the set of optimal solutions over $P$ with respect to the lexicographic objective function $(c_{1}^{\top}x,\dots,c_{i}^{\top}x)$. Let $F^{Q}_{i}$ and $F^{R}_{i}$ be defined analogously. We show by induction on $i$ that $F^{P}_{i}+F^{Q}_{i}=F^{R}_{i}$. For $i=1$, this follows from Lemma 4.4. For $i>1$, observe that $F^{P}_{i}$ is the face in $c_{i}$-direction of $F^{P}_{i-1}$, and analogously for $Q$ and $R$. Hence, we can apply Lemma 4.4 again to complete the induction step. As a consequence, we obtain that $F^{P}_{d}+F^{Q}_{d}=F^{R}_{d}$. Moreover, since $C$ is a basis of $\mathbb{R}^{d}$, it follows that any two optimal solutions with respect to the lexicographic objective function must be equal, implying that $F^{P}_{d}$, $F^{Q}_{d}$, and $F^{R}_{d}$ are singletons. ∎
Proof of Theorem 4.6..
@@
Let $x^{P},x^{Q},x^{R}$ be the unique optimizers as guaranteed by Proposition 4.5. As they fulfill $x^{P}+x^{Q}=x^{R}$, the algorithm correctly returns $x^{P}=x^{R}-x^{Q}$. Furthermore, since $x^{P}$ is optimal with respect to the lexicographic objective, it is in particular optimal with respect to the first component of the objective, which is $c^{\top}x$. Therefore, the algorithm returns an optimal solution for $\max_{x\in P}c^{\top}x$.